# The Final-Stage Bottleneck: A Systematic Dissection of the R-Learner for Network Causal Inference

### v1

@@ 7 Conclusion
Appendix A Experimental Configurations
This appendix contains the full YAML configuration files used to generate the results for each of our main, robustness, semi-synthetic, and control experiments. This is provided for full reproducibility. All experiments were orchestrated by the main script run_experiments.py, which reads one of the following configuration files.
A.1 Configuration: Main Result (BA Graph)
@@
A.2 Configuration: Robustness Check (ER Graph)
@@
A.3 Configuration: Robustness Check (SBM Graph)
@@
A.4 Configuration: Robustness Check (Interaction CATE)
@@
A.5 Configuration: Semi-Synthetic Experiment (Cora)
@@
A.6 Configuration: Negative Control (Local CATE)

### v2

@@ 7 Conclusion
Appendix A A Constructive Example of the Representation Bottleneck
To formally illustrate the existence and nature of the representation bottleneck, we construct a minimal, analytical example. Consider a simple 5-node star graph where node $C$ is the central hub connected to four periphery nodes $A,B,D,E$.
Let each node have a single local feature, defined as:
$X_{A}=1,\quad X_{B}=1,\quad X_{D}=1,\quad X_{E}=2,\quad X_{C}=10$ (6)
Let the network embedding $H_{i}$ for a node $i$ be a function of its neighbors’ features (for simplicity, the neighbor’s feature if degree is 1, or the average of neighbors’ features otherwise). The embeddings are thus:
$H_{A}=H_{B}=H_{D}=H_{E}$ $=X_{C}=10$ (7)
$H_{C}$ $=\frac{1}{4}(X_{A}+X_{B}+X_{D}+X_{E})=\frac{1}{4}(1+1+1+2)=1.25$ (8)
This construction creates two distinct structural roles: the hub (with $H=1.25$) and the periphery (with $H=10$). Now, we define the true CATE $\tau$ to be a non-linear function of this structural embedding, mirroring our main DGP: $\tau_{i}=\sin(H_{i})$. This yields a genuinely heterogeneous, network-dependent CATE:
$\tau_{A}=\tau_{B}=\tau_{D}=\tau_{E}$ $=\sin(10)$ (9)
$\tau_{C}$ $=\sin(1.25)$ (10)
Now, consider a powerful, non-linear but graph-blind final-stage estimator, such as an MLP, whose functional form is restricted to $\hat{\tau}_{i}=f(X_{i})$. To perfectly recover the true CATEs, this model must learn a function $f$ that satisfies the following constraints based on the observable local features $X_{i}$:
$f(1)$ $=\sin(10)\quad(\text{from periphery nodes A, B, D})$
$f(2)$ $=\sin(10)\quad(\text{from periphery node E})$
$f(10)$ $=\sin(1.25)\quad(\text{from hub node C})$
The model is thus forced to map two different inputs, $X=1$ and $X=2$, to the exact same output value, $\sin(10)$. More fundamentally, the input to the function, $X_{i}$, is the wrong information. The true CATE is a function of the structural role (hub vs. periphery), which is captured by $H_{i}$, not by the local feature $X_{i}$.
A graph-aware model, $\hat{\tau}_{i}=g(X_{i},\mathcal{G})$, can trivially learn this function by learning to compute the embedding $H_{i}$ and then approximating $\sin(\cdot)$. A graph-blind model, however, is being asked to learn a function from a variable that has a complex, non-injective relationship with the true causal effect. It is, by construction, theoretically misspecified. This simple construction proves that even a highly flexible non-linear model will fail if it lacks the correct graph-aware inductive bias, thus demonstrating the fundamental nature of the representation bottleneck.
Appendix B Experimental Configurations
This section contains the full YAML configuration files used to generate the results for each of our main, robustness, semi-synthetic, and control experiments. This is provided for full reproducibility. All experiments were orchestrated by the main script run_experiments.py, which reads one of the following configuration files.
B.1 Configuration: Main Result (BA Graph)
@@
B.2 Configuration: Robustness Check (ER Graph)
@@
B.3 Configuration: Robustness Check (SBM Graph)
@@
B.4 Configuration: Robustness Check (Interaction CATE)
@@
B.5 Configuration: Semi-Synthetic Experiment (Cora)
@@
B.6 Configuration: Negative Control (Local CATE)